a 0.05 kg tennis ball approaches a rocket at 25 m/s. if it is in contact with the racket's strings for 0.005 s, then rebounds at 25 m/s, what is the average contact force between the ball and the racket? Ignore the air resistance.


1
Expert's answer
2021-03-23T08:14:22-0400
"F_{avg}\Delta t=m\Delta v,""F_{avg}\Delta t=m(v_f-v_i),""F_{avg}=\dfrac{m(v_f-v_i)}{\Delta t},""F_{avg}=\dfrac{0.05\ kg\cdot(-25\ \dfrac{m}{s}-25\ \dfrac{m}{s})}{0.005\ s}=-500\ N."

The sign minus means that the average contact force between the ball and the racket directed away from the racket.